The potential difference across the heater is 250V. What is the resistance heater?
1
Expert's answer
2018-08-15T09:00:08-0400
The Ohm’s law states
I=V/R
I=q/t
Therefore, the resistance
R=V/I=V/(q/t)=Vt/q=(250 V×60 s)/(240 C)=62.5 Ω